# https://ku-cvlab.github.io/RAIN-GS

### v1

@@ 0.A.3 Ablation on core components details
In Table 4 of Section 6.4, to demonstrate the effectiveness of our progressive Gaussian low-pass filter control strategy, we employed an ablation study utilizing different decreasing functions of convex, linear, and concave to control the Gaussian low-pass filter value $s$. Different from our strategy, where the value $s$ is defined adaptively by image height, width, and the number of Gaussians $N$ at each timestep, the remaining functions were manually defined to achieve $s=300$ at step 0 and $s=0.3$ at about 3,000 steps across all scenes.
@@
As using larger values for the Gaussian low-pass filter enables each Gaussian to model larger regions, this leads to more Gaussians overlapping to each of the tiles which indicates that more Gaussians need to be considered during the alpha blending process. We empirically find that the initial Gaussian low-pass filter value over $s>300$ offers no significant improvement over $s=300$, only making the overall computation inefficient. Based on these findings, we define the max value of the Gaussian low-pass filter as $s=300$. The intuition of our choice of designing convex, linear, and concave functions to reach $s=0.3$ at step 3,000 is based on our analysis that our adaptive Gaussian low-pass filter value reaches $0.3$ between 2,000-3,000 steps. The illustration of different Gaussian low-pass value formulas is shown in Figure 10 and Figure 10 where our formula is adaptively defined, showing different functions for each scene.
@@
Appendix 0.B Proof
0.B.1 Proof on radius of a Gaussian convolved with a low-pass filter
As mentioned in Section 5.2, the 3D Gaussians $G_{i}$ is projected to 2D Gaussians $G_{i}^{\prime}$ in the screen space as follows:
Following the convolution process, 3DGS estimates the projected 2D Gaussian’s area to identify its corresponding screen tiles. This is done by calculating $k$ times the square root of the larger eigenvalue of $(\Sigma_{i}^{\prime}+sI)$, which represents the radius of the approximated circle, and $k$ is the hyperparameter that determines the confidence interval of the 2D Gaussian.
0.B.2 Proof on progressive low-pass filter size
In Section 5.2, we define the value $s$ for our progressive Gaussian low-pass filter control based on the fact that the area of the projected 2D Gaussians is at least $9{\pi}s$. As the area of the projected 2D Gaussian is defined as the circle whose radius is $k$ times the square root of the larger eigenvalue of $(\Sigma_{i}^{\prime}+sI)$, we have to first calculate the eigenvalues of $(\Sigma_{i}^{\prime}+sI)$. If we define the eigenvalues of $\Sigma_{i}$ as $\lambda_{i1},\lambda_{i2}$, since the eigenvalue of $sI$ is $s$, the eigenvalues of $(\Sigma_{i}^{\prime}+sI)$ can be defined as $\lambda_{i1}+s,\lambda_{i2}+s$. This leads to the following proof:
@@ 0.B.2 Proof on progressive low-pass filter size
Appendix 0.C Toy experiments
0.C.0.1 Implementation details.
In this section, we provide a detailed explanation of the 1D regression toy experiment shown in Section 4.2. We design the experiment to learn a target distribution $Y(x)$, where $x$ is in the range $[0,10,000)$, using $N$ 1D Gaussians. The target distribution $Y(x)$ is generated from random 10 1D Gaussian distributions. Each of the $N$ 1D Gaussians includes learnable parameters of $\{\mu_{i},\sigma_{i},w_{i}\}$ and is trained through the L1 loss between the blended distribution from $N$ 1D Gaussians with the weight $w_{i}$ and $Y(x)$ as follows:
Appendix 0.D Additional analysis
0.D.1 Analysis on dense-large-variance (DLV) initialization
Figure 11: Toy experiment learned from different dense initializations visualized in different steps. This figure visualizes the result of our toy experiment predicting the target distribution using a collection of 1D Gaussians, starting from different dense initialization methods. The top row shows the result with dense-small-variance (DSV) initialized Gaussians whereas the bottom row shows the result with dense-large-variance (DLV) initialized Gaussians.
In Section 4.2, we compare the influence of different dense initialization methods through the toy experiment of learning to regress a random 1D signal with multiple 1D Gaussians. Through the toy experiment, we analyzed that while Gaussian learning from wider areas (enabled by the large initial variance) leads to prioritizing low-frequency components, the dense initialization results in instability and convergence issues. As shown in Figure 11, starting from Gaussians with large variance (dense-large-variance (DLV)) leads to prioritizing the learning of low-frequency components, compared to dense-small-variance (DSV) initialization, observed in the coarse signal in the bottom row of step 10.
Table 6: Quantitative comparison on Mip-NeRF360 dataset with various initilize methods. We compare our method with the DSV random initialization method described in the original 3DGS [13] and DLV random initilization method. We report PSNR, SSIM, LPIPS.
  Outdoor Scene
Method bicycle flowers garden stump treehill
@@ 0.D.1 Analysis on dense-large-variance (DLV) initialization
  Indoor Scene
room counter kitchen bonsai average
@@
0.D.2 Analysis on prioritized learning of low-frequency components
In Section 4.1, we find that SfM initialization guides 3D Gaussians with a coarse approximation of the true distribution to robustly learn the remaining high-frequency components. In Section 4.2, we then show that sparse-large-variance (SLV) random initialization successfully guides the Gaussians to prioritize the learning of low-frequency components. Based on these findings, we propose a novel strategy (RAIN-GS) for 3DGS consisting of two main components: 1) sparse-large-variance (SLV) random initialization and 2) progressive Gaussian low-pass filter control. To verify that our strategy successfully guides 3DGS to prioritize the learning of low-frequency components, we provide additional analysis of our strategy through the rendered images achieved from different steps of training.
Figure 13: Images from Figure 12 analyzed in the frequency domain.
As mentioned in Section 4.1, 3DGS (top row) starts from a coarse approximation of the signal, which can be observed by the low-frequency components already being captured in step 100 ((b), top row). This is more evident when compared to the transformed signal of 3DGS (DSV) and Ours at step 100 ((b), middle and bottom row), where the low-frequency components are largely different from the GT signal. Aligned with our analysis of the toy experiment in Section 4.2, DSV random initialization exhibits a tendency towards over-fast convergence on high-frequencies. This can be observed by comparing (b) and (c), where the high-frequency components are quickly learned in (c). This leads to high-frequency artifacts after optimization, as shown in the rendered image (step 30,000, middle row) in Figure 12.
0.D.3 Comparison of computational resources
@@
In this section, we now compare the computational resources of Plenoxels [39], InstantNGP [22], 3DGS [13], 3DGS with DSV random initialization, and 3DGS with our strategy. We evaluate different methods on the Mip-NeRF360 [4], Tanks&Temples [14], and Deep Blending [10] dataset.
0.D.3.1 Training time.
@@
Note that when pre-defined camera poses are available, methods other than 3DGS [13] can bypass the time-consuming Structure-from-Motion (SfM) process. To highlight this overhead, in Table 7 we compare the training time of 3DGS with the SfM process taken into account. For 3DGS, we find that the number of Gaussians in each training step has a direct correlation to the overall training time - more Gaussians lead to longer training time. Our strategy, by utilizing sparse Gaussians in the early stage of training, reduces the training time compared to DSV random initialization.
@@
Table 7: Training time. [^4] Except for the SfM processing time and training times of 3DGS(DSV) and Ours, we copied the values from 3DGS [13]. Note that 3DGS utilized a single A6000 GPU and we utilized a single RTX 3090 GPU for evaluation.
@@
3DGS [^5] The processing time of SfM is denoted with a $\dagger$. 41m33s $+$ (${}^{\dagger}$ 26m54s $+$ (${}^{\dagger}$ 36m2s $+$ (${}^{\dagger}$
3DGS (DSV) 28m22s 18m31s 29m33s
Ours 25m47s 15m50s 24m4s
0.D.3.2 Memory size.
In Table 8, we compare the memory size of different methods. InstantNGP [21], which utilizes a lightweight hash table outperforms all other methods in terms of memory using less than 50MB. As our strategy guides Gaussians to model larger regions in the early stages of training, we find that this leads to an additional advantage of modeling the scene with sparser Gaussians even after sufficient steps of training. This successfully reduces the memory requirements of Ours compared to other 3DGS approaches.
Table 8: Memory size. [^7] Except for 3DGS(DSV) and Ours, we copied the memory size of different models from 3DGS [13].
  Mip-NeRF360 Tanks&Temples Deep Blending
Plenoxels 2.1GB 2.3GB 2.7GB
INGP-Base 13MB 13MB 13MB
INGP-Big 48MB 48MB 48MB
3DGS 734MB 411MB 676MB
3DGS (DSV) 592MB 365MB 582MB
Ours 581MB 359MB 545MB
0.D.4 Limitations
Although RAIN-GS robustly guides 3D Gaussians to model the scene even from randomly initialized point clouds, our novel strategy is not without limitations. As our strategy prioritizes learning the coarse approximation, it sometimes fails to detect the need for further densification to capture high-frequency details. This happens especially for regions where the L1 rendering loss cannot distinguish between a coarse approximation and a high-frequency true distribution. An example of this limitation is shown in Figure 14, where the grass region in the front of the rendered image lacks high-frequency details compared to the ground truth image. However, we find that this limitation also appears when training 3DGS with SfM initialized point clouds, which is also shown to start optimization from a coarse approximation of the true distribution in our main paper. We posit that this limitation mainly comes from the lack of supervision, using the L1 rendering loss as the main signal. Although adding additional supervision from depth or error maps could open up new possibilities to resolve these limitations, we leave this direction as future work.
Figure 14: Visualization of failure cases of learning the coarse approximation first. This figure shows the rendering results of the ’flowers’ scene of the Mip-NeRF360 dataset from Ours and 3DGS. When compared to the GT image shown in (c), the front of the images of (a) and (b) lack high-frequency details.
Appendix 0.E Additional qualitative results
We show additional qualitative results of 3DGS trained with our strategy in Figure 15.
@@ Appendix 0.E Additional qualitative results
Appendix 0.F Additional results of 3DGS in sparse view settings
@@
Table 9: Quantitative comparison on Mip-NeRF360 dataset in sparse-view setting. We compare our method with the random initialization method (DSV) described in the original 3DGS [13]. We report PSNR, SSIM, LPIPS.
  Outdoor Scene
Method bicycle flowers garden stump treehill
@@
3DGS (DSV) 11.787 0.170 0.637 10.948 0.118 0.648 16.752 0.377 0.444 15.239 0.194 0.602 12.675 0.206 0.599
Ours 13.107 0.181 0.626 12.413 0.181 0.584 18.310 0.469 0.416 15.448 0.191 0.594 14.460 0.249 0.580
  Indoor Scene
room counter kitchen bonsai average
@@
3DGS (DSV) 18.864 0.702 0.418 15.311 0.575 0.475 18.181 0.629 0.384 13.246 0.508 0.538 14.778 0.386 0.527
Ours 20.536 0.745 0.404 18.557 0.666 0.432 22.053 0.745 0.332 20.421 0.705 0.422 17.256 0.459 0.488
Table 10: Quantitative comparison on Tanks&Temples and Deep Blending dataset in sparse-view setting. We compare our method with the random initialization method (DSV) described in the original 3DGS [13]. We report PSNR, SSIM, LPIPS.
  Tanks&Temples Deep Blending
Method Truck Train DrJohnson Playroom
[…]
3DGS (DSV) 14.803 0.534 0.434 14.139 0.506 0.478 13.614 0.578 0.595 12.868 0.640 0.579
Ours 16.387 0.591 0.408 14.268 0.489 0.489 17.013 0.644 0.511 16.158 0.693 0.498

### v2

@@ D.2 Ablation study on warm up phase
To demonstrate the effectiveness of our progressive Gaussian low-pass filter control strategy, we employ three different decreasing functions of convex, linear, and concave to control the Gaussian low-pass filter value $s$. Different from our strategy, where the value $s$ is defined adaptively by image height, width, and the number of Gaussians $N$ at each time step, the remaining functions are manually defined to achieve $s=300$ at step 0 and $s=0.3$ at about 3,000 steps across all scenes. The intuition behind this design is based on our analysis that our adaptive Gaussian low-pass filter value reaches $0.3$ between 2,000-3,000 steps. Also, we empirically find that the initial Gaussian low-pass filter value $s>300$ offers no significant improvement, only making the overall computation inefficient. Based on these findings, we define the max value of the Gaussian low-pass filter as $s=300$.
@@
The illustration of different Gaussian low-pass value formulas is shown in Figure 11 and Figure 11 where our formula is adaptively defined, showing different functions for each scene.
Comparison.
In Table 11, we show a more detailed ablation study on the components belonging to the warm up phase. Additional experiments regarding (1) fixing learning rate during the warm up phase (2) the other progressive low-pass filter control functions mentioned above are shown here. Results show that our choice of dynamically adjusting the low-pass filter value based on the number of Gaussians achieves the best performance compared to the others. Also as can be seen in the Table, each of fixing learning rate, SLV initialization and ABE-Split solely contributes to the performance improvement, verifying the effect of our proposed methods. Compared to the original DSV random initialization, we achieve an significant average PSNR improvement over 5dB.
@@
Table 7: Ablation study on warm up phase.
Low-pass filter Init. Fixed learning rate ABE-Split Mip-NeRF360
@@
Constant DSV 22.190 0.704 0.313
Constant SLV 25.675 0.749 0.288
Progressive (Convex) SLV 25.799 0.756 0.286
Progressive (Linear) SLV 25.898 0.754 0.290
Progressive (Concave) SLV 25.974 0.754 0.288
Ours SLV 26.180 0.761 0.280
Ours SLV 26.362 0.791 0.241
Ours SLV 27.228 0.807 0.229
Table 8: Quantitative comparison on Mip-NeRF360 dataset in sparse SfM points setting. We compare our method with 3DGS in sparse SfM points setting, We report PSNR, SSIM and LPIPS.
Method SfM points Outdoor Scene
bicycle flowers garden stump treehill
@@
3DGS cluster 24.542 0.703 0.289 21.044 0.569 0.365 26.883 0.852 0.118 25.774 0.735 0.250 22.102 0.601 0.376
3DGS <10% 24.456 0.700 0.294 21.260 0.580 0.360 26.832 0.850 0.127 26.165 0.744 0.247 22.260 0.612 0.365
3DGS full 25.246 0.771 0.205 21.520 0.605 0.336 27.410 0.868 0.103 26.550 0.775 0.210 22.490 0.638 0.317
Ours cluster 25.211 0.767 0.211 21.704 0.617 0.321 27.166 0.862 0.103 26.816 0.775 0.204 22.589 0.630 0.328
Ours <10% 25.214 0.769 0.208 21.858 0.619 0.320 27.130 0.861 0.107 26.899 0.777 0.202 22.618 0.632 0.327
Method SfM points Indoor Scene Average
room counter kitchen bonsai
@@
3DGS cluster 29.919 0.899 0.255 28.256 0.893 0.227 30.127 0.917 0.140 31.247 0.932 0.220 26.655 0.789 0.249
3DGS <10% 30.795 0.908 0.243 28.523 0.896 0.223 30.120 0.915 0.140 31.625 0.934 0.221 26.893 0.793 0.247
3DGS full 30.632 0.914 0.220 28.700 0.905 0.204 30.317 0.922 0.129 31.980 0.938 0.205 27.205 0.815 0.214
Ours cluster 30.363 0.909 0.233 28.594 0.901 0.211 31.136 0.923 0.133 31.695 0.935 0.217 27.253 0.813 0.218
Ours <10% 30.843 0.912 0.230 28.703 0.902 0.209 31.364 0.922 0.133 31.829 0.938 0.214 27.384 0.815 0.217
D.3 Ablation on sparse SfM initialization
In Table 8, we compare two methods for creating sparse SfM points to reduce the uncertainty of the SfM point cloud and use our coarse-to-fine method with the SfM point cloud: cluster and <10%. For the cluster, we use HDBSCAN [46] to cluster the points based on the location of the SfM point cloud, and use the centroids of the clusters for the sparse SfM initialization. We conduct experiments using the default hyperparameters in scikit-learn [47] for HDBSCAN [46]. For <10%, we create sparse SfM points by using only the top 10% of points utilizing the reprojection error value that comes out with the SfM output. We confirm that our method achieves additional performance improvements by using the SfM points together.
Table 9: Quantitative comparison on Mip-NeRF360 dataset in noisy initial SfM point cloud settings. We compare 3DGS method with different noisy inital SfM point cloud. We report PSNR, SSIM, LPIPS.
Initialization Outdoor Scene
bicycle flowers garden stump treehill
[…]
SfM 25.246 0.771 0.205 21.520 0.605 0.336 27.410 0.868 0.103 26.550 0.775 0.210 22.490 0.638 0.317
SfM + ${\epsilon}$ 24.289 0.692 0.307 21.028 0.565 0.374 26.515 0.846 0.133 26.028 0.743 0.252 21.728 0.607 0.380
SfM+constant 23.619 0.625 0.358 21.139 0.569 0.364 25.663 0.809 0.163 23.382 0.641 0.335 21.989 0.593 0.380
Initialization Indoor Scene Average
room counter kitchen bonsai
[…]
SfM 30.632 0.914 0.220 28.700 0.905 0.204 30.317 0.922 0.129 31.980 0.938 0.205 27.205 0.815 0.214
SfM + ${\epsilon}$ 30.740 0.901 0.260 23.872 0.838 0.285 24.477 0.877 0.175 25.823 0.875 0.265 24.944 0.771 0.270
SfM+constant 30.790 0.899 0.264 28.041 0.875 0.250 30.203 0.914 0.145 31.006 0.924 0.230 26.204 0.761 0.277
D.4 Analysis on the impact of noise on the initial SfM point cloud
In Table 2, we present detailed results to further investigate the ability of the 3DGS optimization scheme to transport Gaussians to the correct 3DGS locations on Mip-NeRF360 datasets. Here, we conduct this experiments by adding random noise $\epsilon\sim\mathcal{N}(0,1)$ and constant systematic noise, whose value equals 2, to the initial SfM points. The results shown in Table 9 prove that 3DGS strongly depends on the initial point. Figure 12 shows initial SfM points and points with noise $\epsilon$. Even with the small amount of noise, 3DGS fails to move to the correct position.
Figure 12: Visualization of SfM points and points with noise $\epsilon$.
Figure 13: Displacements of Gaussians from initial positions
D.5 Analysis on movement of each Gaussian
To observe the movement of the each Gaussian, we measure how far the Gaussian moves from its initial position during the training. An additional parameter is incorporated to record the initial position, ensuring that even when Gaussians are split or cloned, the initial position parameters are retained. Then, the movement is calculated as difference between the final position of each Gaussian after training and its respective initial position.
We conduct analysis on “Truck” scene of Tanks&Temples dataset, comparing the settings of 3DGS with SfM point initialization, 3DGS with random initialization, and our method. The mean, standard deviation, and the top 1% values of the movements is shown in Table 2. Additionally, Figure 13 shows the overall scene from the same camera viewpoint for each experiment to observe the differences in distribution of overall Gaussians between the beginning and 30,000 step. In case such as 3DGS with SfM initialization and random initialization, the positions of the Gaussians does not change significantly. However, our method shows substantial changes in comparison.
D.6 Expanding Scene bound
As discussed in Section 4 and can be seen in supplementary materials Section D.5, Gaussians struggle to move further from the initialized locations. Therefore, in Table 10, we compare the results according to the initial scene bound and our method using the ABE-Split algorithm on the Mip-NeRF360 dataset to see if it is sufficient to initialize with a sufficiently wide scene bound. Extent $\times$ 3 and Extent $\times$ 10 are the initial scene bounds created by multiplying the camera extent given in the dataset by 3 and 10, respectively, and 3 is used in the original 3DGS. As can be seen in Extent x 10, it is confirmed that simply starting with a wide initial scene bound does not yield sufficient performance.
D.7 Detailed results of Oracle scene bound experiments
To find out if the performance degradation of using noisy or random point cloud solely comes from losing the information of where the Gaussians should be located, we conduct a simple experiment of training 3DGS with random point cloud sampled from the oracle scene bounds. Specifically, here we assume the oracle scene bounds as the final scene bound defined by the final positions of the Gaussians, which are trained from point cloud achieved from SfM. Table 11 presents the results for each scene. The experiment reveals that SfM initialization provides additional benefits to 3DGS beyond merely placing the Gaussians in proper locations, as starting from random point cloud even in oracle scene bounds leads to sub-optimal results compared to the one starting from SfM-initialized point cloud.
Table 10: Quantitative comparison on Mip-NeRF360 dataset within different scene bound settings. We compare our method (SLV) with 3DGS using random initialization within different scale bounding box. We report PSNR, SSIM, LPIPS.
Scene bound Outdoor Scene
bicycle flowers garden stump treehill
[…]
Extent×$\times$3 21.034 0.575 0.378 17.815 0.469 0.403 23.217 0.783 0.175 20.745 0.618 0.345 18.986 0.550 0.413
Extent×$\times$10 23.864 0.643 0.354 20.974 0.553 0.381 25.862 0.823 0.151 24.876 0.696 0.300 22.067 0.590 0.394
Ours 25.042 0.747 0.238 21.762 0.616 0.324 26.884 0.854 0.114 26.680 0.768 0.215 22.528 0.621 0.342
Scene bound Indoor Scene Average
room counter kitchen bonsai
[…]
Extent×$\times$3 29.685 0.894 0.265 23.608 0.833 0.276 26.078 0.893 0.161 18.538 0.719 0.401 22.190 0.704 0.313
Extent×$\times$10 29.834 0.889 0.276 27.441 0.863 0.267 30.424 0.913 0.147 30.482 0.917 0.242 26.203 0.765 0.279
Ours 30.809 0.906 0.247 28.529 0.895 0.223 31.270 0.920 0.137 31.547 0.934 0.218 27.228 0.807 0.229
Table 11: Quantitative comparison on Mip-NeRF360 dataset in oracle scene bound settings. We compare 3DGS method with random initialization in paper setting, SfM initialization and random initialization in oracle scene bound. We report PSNR, SSIM, LPIPS.
Initialization Outdoor Scene
bicycle flowers garden stump treehill
[…]
3DGS (Random) 21.034 0.575 0.378 17.815 0.469 0.403 23.217 0.783 0.175 20.745 0.618 0.345 18.986 0.550 0.413
3DGS (SfM) 25.246 0.771 0.205 21.520 0.605 0.336 27.410 0.868 0.103 26.550 0.775 0.210 22.490 0.638 0.317
3DGS (Oracle) 24.227 0.664 0.328 21.353 0.582 0.356 26.791 0.843 0.131 24.634 0.678 0.318 22.219 0.600 0.378
Initialization Indoor Scene Average
room counter kitchen bonsai
[…]
3DGS (Random) 29.685 0.894 0.265 23.608 0.833 0.276 26.078 0.893 0.161 18.538 0.719 0.401 22.190 0.704 0.313
3DGS (SfM) 30.632 0.914 0.220 28.700 0.905 0.204 30.317 0.922 0.129 31.980 0.938 0.205 27.205 0.815 0.214
3DGS (Oracle) 30.192 0.895 0.265 27.812 0.873 0.251 30.797 0.917 0.140 30.303 0.916 0.238 26.481 0.774 0.267
D.8 Comparison of computational resources
@@
In this section, we now compare the computational resources of Plenoxels [35], InstantNGP [25], 3DGS [1], 3DGS with random initialization, and 3DGS with our strategy. We evaluate different methods on the Mip-NeRF360 [21], Tanks&Temples [44], and Deep Blending [45] dataset.
Training time.
@@
Note that when pre-defined camera poses are available, methods other than 3DGS [1] can bypass the time-consuming Structure-from-Motion (SfM) process. To highlight this overhead, in Table 12 we compare the training time of 3DGS with the SfM process taken into account. For 3DGS, we find that the number of Gaussians in each training step has a direct correlation to the overall training time - more Gaussians lead to longer training time.
@@
Table 12: Training time. [^4] Except for the SfM processing time and training times of 3DGS(Random) and Ours, we copied the values from 3DGS [1]. Note that 3DGS utilized a single A6000 GPU and we utilized a single RTX 3090 GPU for evaluation.
@@
3DGS [^5] The processing time of SfM is denoted with a $\dagger$. 41m33s $+$ (19m9s) 26m54s $+$ (19m43s) 36m2s $+$ (6m23s)
3DGS(Random) 28m22s 18m31s 29m33s
Ours 31m34s 15m13s 27m51s
D.9 Limitations
Although RAIN-GS robustly guides 3D Gaussians to model the scene even from randomly initialized point cloud, our novel strategy is not without limitations. As our strategy prioritizes learning the coarse approximation, it sometimes fails to detect the need for further densification to capture high-frequency details. This happens especially for regions where the L1 rendering loss cannot distinguish between a coarse approximation and a high-frequency true distribution. An example of this limitation is shown in Figure 14, where the grass region in the front of the rendered image lacks high-frequency details compared to the ground truth image. However, we find that this limitation also appears when training 3DGS with SfM initialized point cloud, which is also shown to start optimization from a coarse approximation of the true distribution in our main paper. We posit that this limitation mainly comes from the lack of supervision, using the L1 rendering loss as the main signal. Although adding additional supervision from depth or error maps could open up new possibilities to resolve these limitations, we leave this direction as future work.
Figure 14: Visualization of failure cases. This figure shows the rendering results of the “flowers” scene of the Mip-NeRF360 dataset from Ours and 3DGS. When compared to the GT image shown in (c), the front of the images of (a) and (b) lack high-frequency details.
Appendix E Additional qualitative results
We show additional qualitative results in Figure 15.
@@ Appendix E Additional qualitative results
Figure 16: Additional qualitative results with depth.
Appendix F Additional results of 3DGS in sparse view settings
@@
Table 13: Quantitative comparison on Mip-NeRF360 dataset in sparse-view setting. We compare our method with the random initialization method described in the original 3DGS [1]. We report PSNR, SSIM, LPIPS.
Method Outdoor Scene
bicycle flowers garden stump treehill
@@
3DGS(Random) 11.787 0.170 0.637 10.948 0.118 0.648 16.752 0.377 0.444 15.239 0.194 0.602 12.675 0.206 0.599
Ours 12.436 0.157 0.636 11.546 0.120 0.649 17.955 0.463 0.425 14.675 0.142 0.614 12.820 0.191 0.597
Method Indoor Scene Average
room counter kitchen bonsai
@@
3DGS(Random) 18.864 0.702 0.418 15.311 0.575 0.475 18.181 0.629 0.384 13.246 0.508 0.538 14.778 0.386 0.527
Ours 19.804 0.722 0.424 18.201 0.637 0.451 21.474 0.718 0.359 20.664 0.721 0.408 16.620 0.430 0.507